OpenStudy (anonymous):

If f(x)=e^(1/x) then f'(x)=

OpenStudy (precal):

use the chain rule

OpenStudy (anonymous):

\[e^{\frac{1}{x}}\times \frac{d}{dx}[\frac{1}{x}]\] by the chain rule

OpenStudy (anonymous):

I don't get the chain rule

OpenStudy (anonymous):

I got it thanks

OpenStudy (precal):